# Large Language Models for Networking: Workflow, Advances and Challenges

### v1

@@ Abstract
The networking field is characterized by its high complexity and rapid iteration, requiring extensive expertise to accomplish network tasks, ranging from network design, diagnosis, configuration and security. The inherent complexity of these tasks, coupled with the ever-changing landscape of networking technologies and protocols, poses significant hurdles for traditional machine learning-based methods. These methods often struggle to generalize and automate complex tasks in networking, as they require extensive labeled data, domain-specific feature engineering, and frequent retraining to adapt to new scenarios. However, the recent emergence of large language models (LLMs) has sparked a new wave of possibilities in addressing these challenges. LLMs have demonstrated remarkable capabilities in natural language understanding, generation, and reasoning. These models, trained on extensive data, can benefit the networking domain. Some efforts have already explored the application of LLMs in the networking domain and revealed promising results. By reviewing recent advances, we present an abstract workflow to describe the fundamental process involved in applying LLM for Networking. We introduce the highlights of existing works by category and explain in detail how they operate at different stages of the workflow. Furthermore, we delve into the challenges encountered, discuss potential solutions, and outline future research prospects. We hope that this survey will provide insight for researchers and practitioners, promoting the development of this interdisciplinary research field.
@@ I Introduction
Networks, including wide-area internet, data center networks, and satellite networks, are critical infrastructures with distinct design principles and architectures. Due to the complexity and rapid iteration in the field of networking, managing and maintaining these networks involves complex operations that demand rich network expertise. Efficient solutions for network tasks such as network design, network diagnosis, network configuration, and network security are essential for ensuring network stability and scalability. These tasks have garnered significant attention from both academia and industry. Although several machine learning (ML)-based methods have been proposed and shown effectiveness, they often lack generalization and necessitate repeated designs for different scenarios [1].
Recently, there has been significant progress in large language models (LLMs), representing the latest advancement in generative AI. Pretrained on extensive data, these models reveal remarkable capabilities in concept understanding, mathematical reasoning, and tool usage, enabling groundbreaking advancements in various domains, such as chip design, protein structure generation, and robot-embodied intelligence. Inspired by these promising endeavors, some researchers have also explored the application of LLMs in the field of networking and demonstrated encouraging results. In particular, LLMs can benefit the field of networking for the following reasons: (1) LLMs can rapidly learn and grasp network-related knowledge, providing the required expertise for completing various tasks; (2) LLMs can autonomously analyze and reason based on the context of network tasks, offering potential solutions for tackling complex tasks; (3) LLMs can provide a unified natural language interface, making it easier for network operators to execute various tasks.
@@
We then conduct a targeted review (Section III) of notable advancements in networking facilitated by LLMs, predominantly sourced from recent publications. For example, LLMs can assist in understanding protocols and managing network topologies [2, 3], enhance configuration efficiency while minimizing errors [4, 5], identify hidden patterns or anomalies [6, 7], and introduce novel strategies for improving network security [8]. We categorize these advancements into several key fields of networking, including network design, network configuration, network diagnosis, and network security, and elaborate on how they are performed at each stage of the LLMN workflow.
@@ II Basic Workflow for LLMN
Prompt Engineering. Crafting effective prompts is crucial for guiding LLMs to produce desired outputs in networking tasks. Prompts should include descriptions of the network environment, such as network topology, device configurations, and protocol specifications. This helps LLMs understand the task context and generate outputs that align with the network environment. However, due to prompt length limitations, we need to include only closely relevant information, which can be achieved through Retrieval Augmented Generation (RAG) techniques. Additionally, LLM outputs often serve as inputs for other tools or instructions for network device operations, requiring adherence to specific formats. To address these challenges, few-shot learning and Chain-of-thought are proposed to guide LLMs in generalizing and generating accurate results with limited input-output examples, proving to be highly effective. For example, to generate machine-readable grammars for different message types in the RTSP protocols, the prompt includes context like the PLAY client request grammar for RTSP and the GET client request grammar for HTTP [8].
@@
Tool Integration. LLMs excel at reading comprehension and semantic analysis, generating answers based on a pre-established probability distribution in an auto-regressive manner. However, they may struggle with symbolic reasoning and mathematical deduction. To enhance the accuracy of numerical results, crucial for decision-making in networking, one approach is to leverage mathematical tools like Wolfram  [^3] Wolfram: https://www.wolframalpha.com/ or utilize Python code for data analysis and calculation [2, 7, 3]. In addition to general tools, detailed descriptions and usages of domain-specific tools can be provided to enable LLMs to automatically determine and select the appropriate tool based on different task requirements. For example, when it comes to synthesizing router configurations, LLMs alone are not effective. However, by incorporating existing configuration verifiers like Batfish  [^4] Batfish: https://batfish.org/ and Campion  [^5] Campion: https://github.com/atang42/batfish/tree/rm-localize, which offer actionable feedback to address uncertainties and invalidities, LLMs can significantly reduce the need for human intervention [4].
@@ III Overview of Recent Advances
Sharma et al.[2] Design Extraction of FSM from RFC Using Python data structures to represent protocol states and state transitions Greedily engineering seed prompt Off-the-shelf GPT-3.5-turbo ‘pygraphviz’ package /
@@ III-A Network Design
The Network Design task involves detailed planning processes like choosing protocols, allocating bandwidth, and optimizing network topology. The challenge lies in balancing optimal decisions against the backdrop of ever-evolving technology, complex requirements, and limited resources. Traditionally, network engineers have depended on their expertise, manual fine-tuning, and basic tools for network design and optimization. However, LLMs offer a promising avenue by leveraging the extensive knowledge of network professionals and combining it with advanced analytical tools. LLMs can enhance network design by enabling more accurate cost predictions, suggesting tailored strategies for specific network goals, and facilitating the automatic evaluation of various design options.
As a part of initial investigations, several literatures focus on the critical importance of protocol understanding and topology management in network architecture. Sharma et al. [2] introduced PROSPER, a framework leveraging LLMs for the automatic extraction of Finite State Machines (FSMs) from network protocols such as TCP and DCCP, detailed in Requests for Comments (RFCs) by the Internet Engineering Task Force (IETF). This process aids network engineers in deeply understanding protocols and designing networks to meet specific criteria. PROSPER utilizes the Artifact Miner tool to parse diagrams in RFCs, integrating these with the textual content to identify FSM key entities: source state, destination state, and trigger event. These entities, structured as lists of lists or hashmaps of lists, are input into an LLM to produce Python code for FSM diagrams. Experimentally, PROSPER outperforms existing methods, generating FSM diagrams with higher accuracy and fewer errors. Mani et al. [3] have investigated the use of LLMs for code generation in graph analysis and manipulation for topology management. They proposed a structured framework consisting of application wrappers, application prompt generators, code-generation prompt generators, and execution sandboxes. This framework generates user query-based prompts for specific network applications, guides LLMs in producing practical code, and provides a secure execution environment for this code. Their findings reveal that this method significantly improves code quality over direct LLM processing of raw network data, although code quality tends to decline with increasing task complexity. The study suggests that future efforts should aim at enhancing LLMs’ capabilities in managing complex network tasks.
@@ III-B Network Configuration
The evolution of networking and cloud computing technologies has led to the proliferation of network devices, each running various applications to deliver services. Ensuring these devices operate correctly is critically dependent on precise network configuration, a task complicated by the myriad of settings and parameters that demand a deep grasp of networking concepts and operational details. Moreover, the dynamic nature of network conditions and business objectives requires configurations to be regularly updated to align with emerging needs. Traditionally, network configurations have been manually crafted and verified, a process that is not only time-consuming but also prone to errors, making it ill-suited for the dynamic and complex nature of modern networks and their rapidly evolving requirements. Consequently, there is a pressing need for automated solutions that can enhance configuration efficiency, minimize human errors, and swiftly adapt to changes.
@@ III-D Network Security
Mutation-based protocol fuzzing is a security testing technique that assesses protocol implementations by fuzzing seed messages, aiming to uncover potential vulnerabilities. However, the limited diversity of seed messages poses a constraint. To address this limitation, Meng et al. [8] proposed ChatAFL, a novel engine that enhances existing protocol fuzzing tools using LLMs. ChatAFL builds upon AFLNet and introduces machine-readable protocol grammar extraction to enrich seed inputs’ diversity. Moreover, when the fuzzer exhausts exploration of new states, ChatAFL interacts with the LLM to generate client requests that induce server transitions. Experimental results demonstrate that compared to the baseline AFLNet, ChatAFL achieves improved state transition coverage, state coverage, and code coverage, while also identifying more security vulnerabilities.

### v2

@@ Abstract
The networking field is characterized by its high complexity and rapid iteration, requiring extensive expertise to accomplish network tasks, ranging from network design, configuration, diagnosis and security. The inherent complexity of these tasks, coupled with the ever-changing landscape of networking technologies and protocols, poses significant hurdles for traditional machine learning-based methods. These methods often struggle to generalize and automate complex tasks in networking, as they require extensive labeled data, domain-specific feature engineering, and frequent retraining to adapt to new scenarios. However, the recent emergence of large language models (LLMs) has sparked a new wave of possibilities in addressing these challenges. LLMs have demonstrated remarkable capabilities in natural language understanding, generation, and reasoning. These models, trained on extensive data, can benefit the networking domain. Some efforts have already explored the application of LLMs in the networking domain and revealed promising results. By reviewing recent advances, we present an abstract workflow to describe the fundamental process involved in applying LLM for Networking. We introduce the highlights of existing works by category and explain in detail how they operate at different stages of the workflow. Furthermore, we delve into the challenges encountered, discuss potential solutions, and outline future research prospects. We hope that this survey will provide insight for researchers and practitioners, promoting the development of this interdisciplinary research field.
@@ I Introduction
Networks, including wide-area internet, data center networks, and satellite networks, are critical infrastructures with distinct design principles and architectures. Due to the complexity and rapid iteration in the field of networking, managing and maintaining these networks involves complex operations that demand rich network expertise. Efficient solutions for network tasks such as network design, network configuration, network diagnosis, and network security are essential for ensuring network stability and scalability. These tasks have garnered significant attention from both academia and industry. Although several machine learning (ML)-based methods have been proposed and shown effectiveness, they often lack generalization and necessitate repeated designs for different scenarios [1].
Recently, there has been significant progress in large language models (LLMs), representing the latest advancement in generative AI. Pretrained on extensive data, these models reveal remarkable capabilities in concept understanding, mathematical reasoning, and tool usage, enabling groundbreaking advancements in various domains, such as chip design, protein structure generation, and robot-embodied intelligence. Inspired by these promising endeavors, some researchers have also explored the application of LLMs in the field of networking and demonstrated encouraging results. In particular, LLMs offer significant advantages in networking due to their ability to process natural language input and output, eliminating the need for strict data modeling and feature extraction. This reduces task workloads. They can also construct logical chains for complex problem-solving by leveraging domain knowledge, and their transfer learning capability allows them to apply insights from related fields to new network challenges, enhancing their flexibility and applicability.
@@
We then conduct a targeted review (Section III) of notable advancements in networking facilitated by LLMs, predominantly sourced from recent publications. For example, LLMs can assist in designing algorithms and managing network topologies [2, 3], enhance configuration efficiency while minimizing errors [4, 5], identify hidden patterns or anomalies [6, 7], and introduce novel strategies for improving network security [8]. We categorize these advancements into several key fields of networking, including network design, network configuration, network diagnosis, and network security, and elaborate on how they are performed at each stage of the LLMN workflow.
@@ II Basic Workflow for LLMN
Prompt Engineering. Crafting effective prompts is crucial for guiding LLMs to produce desired outputs in networking tasks. Prompts should include descriptions of the network environment, such as network topology, device configurations, and protocol specifications. This helps LLMs understand the task context and generate outputs that align with the network environment. However, due to prompt length limitations, we need to include only closely relevant information, which can be achieved through Retrieval Augmented Generation (RAG) techniques. Additionally, LLM outputs often serve as inputs for other tools or instructions for network device operations, requiring adherence to specific requirements. To address these challenges, few-shot learning and Chain-of-thought are proposed to guide LLMs in generalizing and generating accurate results with limited input-output examples, proving to be highly effective. For example, to generate machine-readable grammars for different message types in the RTSP protocols, the prompt context includes the PLAY client request grammar for RTSP and the GET client request grammar for HTTP [8].
@@
Tool Integration. LLMs excel at reading comprehension and semantic analysis, generating answers based on a pre-established probability distribution in an auto-regressive manner. However, they may struggle with symbolic reasoning and mathematical deduction. To enhance the accuracy of numerical results, crucial for decision-making in networking, one approach is to leverage mathematical tools like Wolfram  [^3] Wolfram: https://www.wolframalpha.com/ or utilize Python code for data analysis and calculation [7, 3]. In addition to general tools, detailed descriptions and usages of domain-specific tools can be provided to enable LLMs to automatically determine and select the appropriate tool based on different task requirements. For example, when it comes to synthesizing router configurations, LLMs alone are not effective. However, by incorporating existing configuration verifiers like Batfish  [^4] Batfish: https://batfish.org/ and Campion  [^5] Campion: https://github.com/atang42/batfish/tree/rm-localize, which offer actionable feedback to address uncertainties and invalidities, LLMs can significantly reduce the need for human intervention [4].
@@ III Overview of Recent Advances
He et al. [2] Design Designing Adaptive Bitrate Algorithms Rewrite Pensieve’s original code of state creation and network architectures into a function respectively CoT & Prompt Template Off-the-shelf GPT-3.5 & GPT-4 / Evaluation of various candidate algorithms in a network simulator
@@
Wang et al. [12] Providing in-depth analysis and customized mitigation recommendations for DDoS attack Converting raw traffic into LLM-understandable texts Prompt template Off-the-shelf GPT-4 / /
@@ III-A Network Design
The Network Design task involves detailed planning processes like protocol selection, bandwidth allocation, and network topology optimization. The challenge lies in making optimal decisions amidst ever-evolving technology, complex requirements, and limited resources. Traditionally, network engineers have depended on their expertise, manual fine-tuning, and basic tools for network design and optimization. However, LLMs offer a promising avenue by combining extensive networking knowledge with advanced reasoning and generation capabilities. LLMs can enhance network design by proposing tailored strategies for specific network goals and facilitating the automatic evaluation of various design options.
As a part of initial investigations, several pieces of literature focus on the critical importance of data transmission and topology management in network architecture. He et al. [2] proposed leveraging LLM to develop ABR (Adaptive Bitrate) algorithms, a pivotal component for dynamically adjusting video quality in streaming media delivery. Given the challenge of directly crafting high-quality algorithms with LLM, they incorporated the classic Pensieve algorithm as part of the input prompts. Their objective was to ask LLM to enhance the original algorithm and generate a range of candidate algorithms. As not all these algorithms exhibited significant enhancements, evaluating these algorithms in a network simulator was imperative to identify the most effective one. Experimental findings demonstrate LLM’s capability to produce outstanding algorithms tailored to diverse network conditions. Mani et al. [3] have investigated the use of LLMs for code generation in graph analysis and manipulation for topology management. They proposed a structured framework consisting of application wrappers, application prompt generators, code-generation prompt generators, and execution sandboxes. This framework generates user query-based prompts for specific network applications, guides LLMs in producing practical code, and provides a secure execution environment for this code. Their findings reveal that this method significantly improves code quality over direct LLM processing of raw network data, although code quality tends to decline with increasing task complexity. The study suggests that future efforts should aim at enhancing LLMs’ capabilities in managing complex network tasks.
@@ III-B Network Configuration
The evolution of networking and cloud computing technologies has led to the proliferation of network devices, each running various applications to deliver services. Ensuring these devices operate correctly is critically dependent on precise network configuration, a task complicated by the myriad of settings and parameters that demand a deep grasp of networking concepts and operational details. Moreover, the dynamic nature of network conditions and business objectives requires configurations to be regularly updated to align with emerging needs. Traditionally, network configurations have been manually crafted and verified, a process that is not only time-consuming but also prone to errors, making it ill-suited for the dynamic and complex nature of modern networks and their rapidly evolving requirements.
@@ III-D Network Security
Mutation-based protocol fuzzing is a security testing technique that assesses protocol implementations by fuzzing seed messages, aiming to uncover potential vulnerabilities. However, the limited diversity of seed messages poses a constraint. To address this limitation, Meng et al. [8] proposed ChatAFL, a novel engine that enhances existing protocol fuzzing tools using LLMs. ChatAFL builds upon AFLNet and introduces machine-readable protocol grammar extraction to enrich seed inputs’ diversity. Moreover, when the fuzzer exhausts exploration of new states, ChatAFL interacts with the LLM to generate client requests that induce server transitions. Experimental results demonstrate that compared to the baseline AFLNet, ChatAFL achieves improved state transition coverage, state coverage, and code coverage, while also identifying more security vulnerabilities. Considering existing AI-driven methods for DDoS defense have certain limitations in practical application, including lack of explainability and lack of mitigation instructions, Wang et al. [12] have proposed a LLM-based DDoS defense framework named ShieldGPT to solve these shortcomings. ShieldGPT comprises four modules: attack detection, traffic representation, domain-knowledge injection, and role representation. It transforms raw traffic data into LLM-friendly text, maintaining essential features. By integrating this data with domain knowledge using prompt templates, ShieldGPT enables the LLM to elucidate attack behaviors and propose specific mitigation actions. Preliminary experiments demonstrate that ShieldGP can provide in-depth attack analysis and customized mitigation recommendations, laying the foundation for establishing a truly autonomous DDoS defense system.